OpenStudy (anonymous):

A bag has four green marbles three red marbles and three yellow Marbles what is the probability that you pick a red marble, do not replace it and pick another red marble

OpenStudy (anonymous):

Total marbles = 4 + 3 + 3 = 10 Chance of picking a red marble on first pick : 3/10 Chance of picking a red marble on second pick: 2/9 Therefore probability = 3/10 * 2/9 = 6/90 = 2/30 = 1/15 b) 1/15

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@Thebigjg99

OpenStudy (anonymous):

Thanks